# Partial Rewriting and Value Interpretation of Logically Constrained Terms

### v1

@@ 4 Interpretations of Partial and Most General Constrained Reductions
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be an existentially constrained term. A term $t$ is said to be an instance of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ if there exists an $X$-valued substitution $\sigma$ such that $\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$ and $t=s\sigma$. We denote the set of all instances of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ by $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$, i.e.,
[…]
Note that by unfolding the definition of $\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$, we have $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle=\{s\sigma\mid\sigma(X\cup\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi))\subseteq\mathcal{V}\mathsf{al},\vDash_{\mathcal{M}}(\exists\vec{x}.\ \varphi)\sigma\}$.
@@
We have $\langle\!\langle\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x,z)\penalty 10000\ [\hskip 0.85358pt\exists y.\ x=y\times 2\hskip 0.85358pt]\rangle\!\rangle=\{\mathsf{f}(n,w)\mid n\in\mathbb{Z}/2,w\in\mathcal{T}(\mathcal{F},\mathcal{V})\}$.
@@
Proof 4.5.
The following properties, which directly follow from the definition, are immediate corollaries of Lemma˜4.4:
@@
The interpretation of rewrite rules is defined analogously, by considering synchronized instances of both sides of each rule.
@@
Let $\rho:\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ be a left-linear constrained rewrite rule. We define the interpretation $\langle\!\langle\rho\rangle\!\rangle$ of $\rho$ as: $\langle\!\langle\rho\rangle\!\rangle=\{\langle\ell\sigma,r\sigma\rangle\mid\mathcal{D}\mathsf{om}(\sigma)=\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)=Z\cap\mathcal{V}\mathsf{ar}(\ell,r),\sigma\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\}$, where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$. For an LCTRS $\mathcal{R}$, we define $\langle\!\langle\mathcal{R}\rangle\!\rangle=\bigcup_{\rho\in\mathcal{R}}\langle\!\langle\rho\rangle\!\rangle$.
@@
Let $\rho:\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x,y)\rightarrow y\penalty 10000\ [\hskip 0.85358pt0\geqslant x\hskip 0.85358pt]$ be a constrained rewrite rule. Then $\langle\!\langle\rho\rangle\!\rangle=\{\mathsf{f}(n,y)\to y\mid 0\geqslant n\in\mathbb{Z}\}$.
Note that, in contrast to the interpretation of existentially constrained terms, non-logical variables of rewrite rules are not instantiated in the interpretation of rewrite rules. We note also that, unlike the interpretation of existentially constrained terms, the non-logical variables in rewrite rules are not instantiated in the interpretation of rewrite rules.

### v2

@@ 4 Interpretations of Partial and Most General Constrained Reductions
Let $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ be an existentially constrained term. A term $t$ is said to be an instance of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ if there exists an $X$-valued substitution $\sigma$ such that $\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi$ and $t=s\sigma$. We denote the set of all instances of $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]$ by $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$, i.e., $\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle=\{\hskip 0.85358pts\sigma\mid\sigma(X)\subseteq\mathcal{V}\mathsf{al},\sigma\vDash_{\mathcal{M}}\exists\vec{x}.\ \varphi\hskip 0.85358pt\}$.
@@
We have $\langle\!\langle\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x,z)\penalty 10000\ [\hskip 0.85358pt\exists y.\ x=y\times 2\hskip 0.85358pt]\rangle\!\rangle=\{\hskip 0.85358pt\mathsf{f}(n,w)\mid n\in\mathbb{Z}/2,w\in\mathcal{T}(\mathcal{F},\mathcal{V})\hskip 0.85358pt\}$; thus, e.g., $\mathsf{f}(0,\mathsf{f}(x,y))\in\langle\!\langle\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x,z)\penalty 10000\ [\hskip 0.85358pt\exists y.\ x=y\times 2\hskip 0.85358pt]\rangle\!\rangle$.
@@
The following properties, which also directly follow from the definition of subsumption and equivalence, are immediate corollaries of Lemma˜4.4:
@@
The interpretation of rewrite rules needs to synchronize instantiations of both sides of each rule.
@@
Let $\rho:\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ be a left-linear constrained rewrite rule. We define the interpretation $\langle\!\langle\rho\rangle\!\rangle$ of $\rho$ as: $\langle\!\langle\rho\rangle\!\rangle=\{\hskip 0.85358pt\langle\ell\sigma,r\sigma\rangle\mid\mathcal{D}\mathsf{om}(\sigma)=\mathcal{V}\mathcal{D}\mathsf{om}(\sigma)=Z\cap\mathcal{V}\mathsf{ar}(\ell,r),\sigma\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\hskip 0.85358pt\}$, where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$. For an LCTRS $\mathcal{R}$, we define $\langle\!\langle\mathcal{R}\rangle\!\rangle=\bigcup_{\rho\in\mathcal{R}}\langle\!\langle\rho\rangle\!\rangle$.
@@
Let $\rho:\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{f}(x,y)\rightarrow y\penalty 10000\ [\hskip 0.85358pt0\geqslant x\hskip 0.85358pt]$ be a constrained rewrite rule. Then $\langle\!\langle\rho\rangle\!\rangle=\{\hskip 0.85358pt\mathsf{f}(n,y)\to y\mid 0\geqslant n\in\mathbb{Z}\hskip 0.85358pt\}$.
We note that, unlike the interpretation of existentially constrained terms, the non-logical variables in rewrite rules are not instantiated in the interpretation of rewrite rules. We define it this way because variable instantiation is built into the rewrite steps.